OpenStudy (anonymous):

How do you write a recursive and explicit formula for the sequence 15, 22, 29, 36, 43, .... ? I hope someone can explain this to me!

OpenStudy (thomas5267):

\[ A_n=A_{n-1}+7 \] Since the last term is 7 more than then current term.

OpenStudy (michele_laino):

hint: we can write this: \[\Large \begin{gathered} {a_2} = 22 = 15 + 1 \cdot 7 = {a_1} + \left( {2 - 1} \right) \cdot 7 \hfill \\ \hfill \\ {a_3} = 29 = 22 + 7 = 15 + 2 \cdot 7 = {a_1} + \left( {3 - 1} \right) \cdot 7 \hfill \\ \hfill \\ {a_4} = 36 = 29 + 7 = 15 + 3 \cdot 7 = {a_1} + \left( {4 - 1} \right) \cdot 7 \hfill \\ \hfill \\ {a_5} = 43 = 36 + 7 = 15 + 4 \cdot 7 = {a_1} + \left( {5 - 1} \right) \cdot 7 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I see, then the \[a_n-1 \] represents the preceding term. How do you write an explicit formula for the sequence?

OpenStudy (anonymous):

Nvm, I think I'm good thank you